Start learning →

In 3 trials with success probability 0.3, find P(exactly 2 successes).

A0.063
B0.027
C0.189
D0.441
Answer & Solution
Correct answer: C. 0.189
P = 3C2 (0.3)^2 (0.7) = 3 x 0.09 x 0.7 = 0.189.
Practise this question with Quill — MHT-CET →

Keep practising Binomial Distribution

Browse all Binomial Distribution questions for MHT-CET, or explore more Mathematics topics.

Start a Binomial Distribution lesson →
Related questions